Using the definition, show' that the sequence [1/√n]neN is Cauchy.

1
Expert's answer
2022-03-30T11:30:01-0400

ANSWER.

By the definition, a sequence "\left \{a _{n} \right \}" is called a Cauchy sequence if for any given "\varepsilon>0" there exists "n_{\varepsilon}\in N" such that

"n,m>n_{\varepsilon }\Rightarrow\left | a_{n} -a_{m}\right |<\varepsilon" .

Note, that if "m \neq n" , then "m>n" (or vice versa). Let "m>n" , then "m=n+(m-n)" Denoting "p=m-n", we get the equivalent definition: "\left \{a _{n} \right \}" is called a Cauchy

sequence if for any given "\varepsilon>0" there exists "n_{\varepsilon}\in N" such that

"\left | a_{n} -a_{n+p}\right |<\varepsilon" for all "n >n_{\varepsilon }" and "p\in N" .

Since "\sqrt{n+p}>\sqrt{n}" for all "n,p\in N" ,then

"\left | a_{n}-a_{n+p} \right |=\frac{1}{\sqrt{n }}- \frac{1}{\sqrt{n+p}}<\frac{1}{\sqrt{n }}" .

So, if "\varepsilon>0" and "n_{\varepsilon}=""\left [ \left ( \frac{1}{\varepsilon} \right )^{2} \right ]" (integer part of "\left ( \frac{1}{\varepsilon} \right )^{2}" ) , then

"n>n_{\varepsilon}>\left ( \frac{1}{\varepsilon} \right )^{2} \Rightarrow \sqrt{n}>\sqrt{n_{\varepsilon}} >\left ( \frac{1}{\varepsilon} \right ) \Rightarrow0< \frac{1}{\sqrt{n}}<\varepsilon" ,

Hense , "0<\frac{1}{\sqrt{n }}- \frac{1}{\sqrt{n+p}}<\varepsilon" "(n>n_{\varepsilon}, p\geq1)" .

Therefore , the sequence "a_{n}=\frac{1}{\sqrt{n}}" is Cauchy sequence